### v2

# Dynamic Tool Dependency Retrieval for Efficient Function Calling

Abstract
Function calling agents powered by Large Language Models (LLMs) select external tools to automate complex tasks. On-device agents typically use a retrieval module to select relevant tools, improving performance and reducing context length. However, existing retrieval methods rely on static and limited inputs, failing to capture multi-step tool dependencies and evolving task context. This limitation often introduces irrelevant tools that mislead the agent, degrading efficiency and accuracy. We propose Dynamic Tool Dependency Retrieval (DTDR), a lightweight retrieval method that conditions on both the initial query and the evolving execution context. DTDR models tool dependencies from function calling demonstrations, enabling adaptive retrieval as plans unfold. We benchmark DTDR against state-of-the-art retrieval methods across multiple datasets and LLM backbones, evaluating retrieval precision, downstream task accuracy, and computational efficiency. Additionally, we explore strategies to integrate retrieved tools into prompts. Our results show that dynamic tool retrieval improves function calling success rates between $23\%$ and $104\%$ compared to state-of-the-art static retrievers.
11footnotetext: Work done during an internship at Qualcomm AI Research, Amsterdam. Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
1 Introduction
Figure 1: Dynamic Tool Dependency Retrieval. Given demonstration data for a set of tools, previous work retrieve tools based on either a) the natural language query (highlighted in orange) or b) the latest executed tool call in the plan (the red triangle, highlighted in teal). We instead propose a retrieval method which is dynamically conditioned on both the query and the current history of tool calls (blue square, then red triangle). This dynamic method allows to retrieve only tools that are strictly relevant to both query (orange) and tool context (teal).
Large language models (LLMs) augmented with tool use (a.k.a., function calling) have rapidly evolved from early neuro-symbolic systems to agentic frameworks that can plan, select, and invoke external Application Programming Interfaces (APIs) (Yao et al. 2023; Schick et al. 2023; Patil et al. 2024; Patel, Jagmohan, and Vempaty 2025). Despite this progress, deploying tool-augmented LLMs on-device remains challenging due to two key constraints: (i) efficiency under strict memory and latency budgets, and (ii) effectiveness across large and heterogeneous tool sets.
Therefore, prior work have proposed using tool retrieval modules to encode only relevant tools into the prompt of the function calling agent (Qin et al. 2023; Braunschweiler, Doddipatla, and Zorila 2025; Paramanayakam et al. 2025). Doing so makes it easier for the agent to identify the correct function, reducing unnecessary calls, and enhancing both accuracy and prompt efficiency. However, a major challenge is determining what information should guide the tool retrieval. Some methods rely solely on semantic similarity between the query and tool descriptions (Gao et al. 2025; Paramanayakam et al. 2025), which can overlook contextual relevance of selected tools. Others leverage static tool-dependency graphs built from demonstration trajectories (Liu et al. 2024a), but these approaches risk retrieving tools irrelevant to the query or being biased toward repeated calls. For retrieval to provide the most informed context, it must adapt to both the current task and the ongoing trajectory, while remaining lightweight enough to satisfy on-device constraints. This setting raises the question: can such specificity be achieved with a low-resource approach? To answer this, we introduce Dynamic Tool Dependency Retrieval (DTDR), a lightweight retrieval component that, given a user query and partial plan, identifies a small set of relevant tools and their dependency relations (see Figure 1). Our main contributions are as follows:
  1. 1.
    Lightweight Tool-Dependency Retrieval method. We formulate Dynamic Tool Dependency Retrieval (DTDR), a dependency-aware tool retrieval framework that conditions on both the user query and the evolving tool plan to recover a minimal, task-specific dependency subgraph. Through comprehensive analysis we provide strong evidence that methods without query- and history-awareness are unable to solve the tool retrieval task, so history-aware methods should be adopted.
  2. 2.
[…]
  3. 3.
    Analysis with multiple Prompt Encoding strategies. We use a prompt-efficient in-context learning (ICL) representation that conditions the LLM only on the minimal subgraph of tool dependencies retrieved. We benchmark multiple ICL encoding strategies, identifying weighted Hard Masking as the best contender, but also investigating when and why other approaches should be preferred based on different factors (model scale, dataset statistics, tool retrieval accuracy).
Method Tool Deps. Aware Tool Desc. Free Query Aware Multi-Step History Aware Small model
BM25 (Robertson, Zaragoza et al. 2009)
ToolGraph Retriever (Gao et al. 2025)
Less-is-More Lv1 (Paramanayakam et al. 2025)
ToolNet (Liu et al. 2024a)
TinyAgent (Erdogan et al. 2024)
Toolformer (Schick et al. 2023)
DTDR (Ours)
Table 1: Comparison of tool retrieval methods used in prior work. Our work is the only one that satisfies all 55 desired conditions.
2 Related Work
Recent LLMs demonstrate impressive tool-usage capabilities, with cloud-based models such as GPT-5 (OpenAI 2025), Claude 4 (Anthropic 2025), and GLM-4.5 (Zeng et al. 2025) leading by a wide margin on function-call benchmarks like BFCL V4 (Patil et al. 2025) and $\tau^{2}$-Bench (Barres et al. 2025). Recent work has explored fine-tuning LLMs on large datasets to build general purposes function calling models (Cheng-Jie Ji et al. 2024). These model often fail in real-world scenarios due to poor function selection, misinterpretation of user intent, or under realistic data perturbations (Dang et al. 2025; Rabinovich and Anaby-Tavor 2025). Alternatively, ICL strategies for tool learning can involve including tool descriptions (Shen et al. 2023, 2024; Patel, Jagmohan, and Vempaty 2025) or example trajectories (Paranjape et al. 2023; Sarukkai, Xie, and Fatahalian 2025) within the model prompt. To handle hundreds of functions, recent work uses retrieval-augmented generation to sub-select tools that are relevant to the task (Qin et al. 2023; Braunschweiler, Doddipatla, and Zorila 2025; Paramanayakam et al. 2025). Most prior work retrieve relevant tools based on the query and tool descriptions (Braunschweiler, Doddipatla, and Zorila 2025; Paramanayakam et al. 2025; Paranjape et al. 2023). (Liu et al. 2024b) and (Ding et al. 2025) assume a known tool dependency graph and use it to aid the LLM with selecting relevant functions. As tool dependencies are often unknown, others propose learning them via demonstrations (Paranjape et al. 2023; Chen et al. 2025; Liu et al. 2024a; Qin et al. 2023; Patil et al. 2024; Erdogan et al. 2024). Unlike prior retrievers that rely solely on the user query or static tool-dependency graphs, we propose a dynamic tool retrieval approach that conditions on both the current query and the sequence of previously invoked tools. A more extensive discussion of related works can be found in Appendix A. In Table 1 in the Appendix, we summarize related works that most closely align with our method. We compare against $5$ different categories. 1) Tool Dependency Aware: retrievers that utilize tool dependencies are better informed to solve multi-step tasks (Gao et al. 2025); 2) Tool Description Free: function documentation may not be available or consistent across large sets of tools (Patel, Jagmohan, and Vempaty 2025); 3) Query Aware: retrieval methods should retrieve tools relevant to the specific task, i.e. email tasks will more likely need getting an email address than opening a document, 4) Multi-Step History Aware: retrievers that take into account multi-step history, not just current step, avoid bias towards frequently called functions, and 5) Small Model: On-device agents have stricter memory and latency requirements that the retrieval method needs to satisfy. Our method is the only one that satisfies all $5$ categories.
3 Problem Formulation
Retrieval-Augmented Function Selection.
Consider the function selection agent $\pi$, consisting of a frozen language model (LM). Let $\mathcal{F}$ be the collection of function names that the agent can choose from to solve a task described through a natural language query $q\in Q$, such as “Reply to my latest email from Willem.”. Let $f_{q,0:t-1}=\left[f_{0},f_{1},...,f_{t-1}\right]$ be the list of previous functions predicted for $q$ by $\pi$ up until timestep $t$. Typically, $q$, $f_{q,0:t-1}$ and $\mathcal{F}$ are encoded in the LM prompt, $p$, which is input to the agent to sample a function $f_{t}\sim\pi(\cdot|p(q,f_{q,0:t-1},\mathcal{F}))$. A sampled function $f_{t}$ is correct if $f_{t}\in\mathcal{F}_{q,t}^{*}$, where $\mathcal{F}_{q,t}^{*}$ is the set of correct functions for the given query $q$ at timestep $t$. To raise the probability of $f_{t}\in\mathcal{F}_{q,t}^{*}$, prior work encodes into the prompt a subset of $\mathcal{F}$ (Qin et al. 2023; Braunschweiler, Doddipatla, and Zorila 2025). This subset is retrieved via some module $\omega(\cdot)$, a function that outputs a subset of function names $\mathcal{F}_{t}\subseteq\mathcal{F}$. The set of inputs to the $\omega$ is dependent on the specific retrieval method. Once again for notation simplicity, we drop $q$ for $f_{0:t}:=f_{q,0:t},\,\mathcal{F}_{t}^{*}:=\mathcal{F}_{q,t}^{*}$. Let $T$ be the maximum trajectory length for any $q\in Q$, let $\Omega$ be the set of all possible $\omega$ retriever functions, and let $\mathbf{1}$ be a scalar indicator function. Thus, our function selection problem is a prompt optimization objective:
$\max_{\omega\in\Omega}\sum_{q\in Q}\sum_{t=0}^{T}\mathbf{E}_{f_{t}\sim\pi(\cdot|p(q,f_{0:t-1},\mathcal{F}_{t}))}[\mathbf{1}_{f_{t}\in\mathcal{F}_{t}^{*}}]$ (1)
This optimization problem provides mathematically grounds the success of the function calling agent $\pi$ to how well the retrieved set $\mathcal{F}_{t}$ aligns with $\mathcal{F}_{t}^{*}$. Specifically, we consider an optimal retrieval module $\omega^{*}$ as one such that it retrieves a non-empty subset of the ground-truth set: $\{\}\subset\omega^{*}(\cdot)=\mathcal{F}_{t}\subseteq\mathcal{F}_{t}^{*}$. Note that we focus on optimizing the individual function selection step for more fine-grained downstream evaluation of the retrieval module, so this objective function is less sparse than ones for multi-step, function-calling tasks (Patel, Jagmohan, and Vempaty 2025). We discuss in Section 6 end-to-end results with trajectory-level evaluation.
Limitation of Prior Work: Suboptimal Sets of Inputs. Some zero-shot retrieval methods (Paramanayakam et al. 2025; Gao et al. 2025) implement $\omega$ as the cosine similarity between embeddings of the query (or a transformation of it) and embeddings of each tool description. Therefore, let $f_{desc}$ be the description of $f$, and letting $\mathcal{F}_{desc}=\{f_{desc}\mid\forall f\in\mathcal{F}\}$, the retriever module $\omega$ becomes a function of $q$ and $\mathcal{F}_{desc}$: $\omega\left(q,\mathcal{F}_{desc}\right)$. Furthermore, the retrieved set $\mathcal{F}_{t}$ is used throughout the entire trajectory, constant throughout time steps $t\in[0,T]$. Relying solely on static tool descriptions often overlook the dynamic nature of tool usage in context. While semantic similarity between queries and tool descriptions may capture surface-level relevance (Lin et al. 2024), it fails to account for how tools are actually sequenced in real tasks. Other work such as (Liu et al. 2024a) utilize collected demonstration trajectory data $D$ of function calling tasks, obtained by tool providers (Paranjape et al. 2023), historical tool trajectories (Chen et al. 2025) or generated (Erdogan et al. 2024; Gao et al. 2025; Patil et al. 2024). With $D$, they compute empirical next-function probabilities based on the latest function $f_{t-1}$ to obtain $\mathcal{F}_{t}$. However, this limited context does not model multi-step tool dependencies well. For example, if sending the same email to multiple people, “get_email_address” must be called consecutively multiple times. Demonstrations that contain function trajectories such as these will cause a high value of $P(\text{{``get\_email\_address''}}|\text{{``get\_email\_address''}})$, thus biasing the agent to repeatedly call it more times than necessary. Furthermore, computing probabilities based on the entire demonstration data across various queries underfits tool-use patterns. The $P(\text{{``get\_email\_address''}}|\text{{``get\_email\_address''}})$ value suitable in tasks with multiple recipients can be misleading in tasks with a single recipient. Without the task-specific information, the context of previous functions may be a misleading prior for $\pi$.
In this work, we aim for a lightweight retrieval mechanism to obtain tool-dependency priors that capture both task-specific and context-specific information for on-device function calling.
4 Proposed Approach: Dynamic Tool Dependency Retrieval
Figure 2: System diagram for DTDR. On the left, the user query and tool history are input to DTDR to retrieve the most likely next tools. The LLM $\pi$ selects the next tool among this set. On the right, we show the two alternative instantiations for the retriever: a) DTDR-C, based on a clustering step to retrieve an explicit graph of tool dependencies; and b) DTDR-L, based on a learned linear classifier implicitly modeling tool dependencies. Both systems are conditioned on user query and full tool history.
To capture appropriate tool-dependencies for Retrieval-Augmented Function Calling, we introduce Dynamic Tool Dependency Retrieval (DTDR). The high-level idea of DTDR is simple: the retrieval module $\omega$ should be based on both the task $q$ and the trajectory of previous function calls $f_{0:t-1}$; thus, $\omega(q,f_{0:t-1})$.
Figure 2 (left) gives an overview of our approach. The test query $q$ and the trajectory of function calls $f_{0:t-1}$ are input to DTDR, which predicts a set of tools $\mathcal{F}_{t}:=\omega(q,f_{0:t-1})$, assigning a probability to each. We then encode the retrieved dependencies into the prompt $p(q,f_{0:t-1},\mathcal{F}_{t})$ by hard masking the full set of tools $\mathcal{F}$, while also providing the probabilities for the retrieved tools. In the Experimental Setup Section, we elaborate on other ICL methods we investigated to embed $\mathcal{F}_{t}$ into the prompt. The encoded prompt is then processed by the LLM agent to sample the next function $f_{t}\sim\pi(\cdot|p(q,f_{q,0:t-1},\mathcal{F}))$. For end-to-end evaluation, this continues in an iterative approach starting from the empty function call plan $f_{\emptyset}=\left[\right]$, and ending after predicting the “end-of-plan” function $f_{t}=\text{{``end''}}$.
4.1 Variants of DTDR
We propose two lightweight variants for DTDR: one supervised gradient-based method and one unsupervised clustering-based method. Firstly, proposing two allows us to validate whether our claims on dynamic tool retrieval are consistent across different categories of methods. Secondly, it enables us to directly compare against previous work belonging to these two categories. Lastly, depending on software and system constraints, either variant might be preferred for a real on-device scenario. Both proposed variants utilize demonstration data $\mathcal{D}$.
Dynamic Tool Dependency Retrieval-Clustering (DTDR-C).
This variant (see Figure 2, top-right) is based on a clustering component and a graph-traversal component. The clustering component maps a test query to a tool dependency graph which is relevant for the specific task. The graph-traversal component traverses the graph based on the history of tool calls, and determines the most likely next tools for the plan. Formally, we embed demonstration queries $Q_{\mathcal{D}}$ with a pretrained embedding model $\theta$ and fit a $K$-Means clustering model $C$ on the embeddings. For each cluster $k$ with $0\leq k<K$, we consider the set of demonstrations assigned to that cluster $D_{k}=\{d\mid q_{d}\in Q_{D},\ C(\theta(q_{d}))=k\}$. We build a weighted tool dependency graph $G_{k}$ describing the next-tool dependency probabilities given a sequence of tools. In particular, $G_{k}(f_{0:t-1})$ describes the set of functions which follow the history $f_{0:t-1}$ in the demonstration data $D_{k}$, as well as their probabilities. Details on constructing the tool dependency graph $G$ are included in Appendix  C. Given a test query $q$ and history $f_{0:t-1}$, the next-tool prediction in DTDR-C can be obtained with $\mathcal{F}_{t}=\omega(q,f_{0:t-1})=G_{C(\theta(q))}(f_{0:t-1})$, or, with simpler notation, $\mathcal{F}_{t}=G_{C,q}(f_{0:t-1})$. The number of learned parameters in the $K$-means model is $e_{\theta}*K$, only depending on the output size of the embedding model $e_{\theta}$ and the number of clusters $K$.
Fun. Sel. Acc. [%] ($\uparrow$)
Retrieval Metrics
Retrieval Method   MRR ($\uparrow$)   F1 ($\uparrow$)
TinyAgent Random Guess 05.9 0.26 0.12
BM-25 (Robertson, Zaragoza et al. 2009) 23.1 0.35 0.20
QTS (Vanilla) 15.8 0.26 0.14
QTS (Gao et al. 2025) 21.5 0.18 0.09
QTS (Paramanayakam et al. 2025) 23.7 0.36 0.19
DR (Liu et al. 2024a) 30.7 0.70 0.49
Dynamic DR (DTDR-C) (Ours) 43.3 0.78 0.56
LR (Erdogan et al. 2024) 25.6 0.53 0.39
Dynamic LR (DTDR-L) (Ours) 65.1 0.93 0.55
Taskbench-HF Random Guess 04.2 0.16 0.05
BM-25 (Robertson, Zaragoza et al. 2009) 08.3 0.18 0.05
QTS (Vanilla) 07.9 0.21 0.08
QTS (Gao et al. 2025) 14.6 0.37 0.27
QTS (Paramanayakam et al. 2025) 12.9 0.30 0.18
DR (Liu et al. 2024a) 17.1 0.46 0.33
Dynamic DR (DTDR-C) (Ours) 27.5 0.64 0.55
LR (Erdogan et al. 2024) 20.5 0.53 0.32
Dynamic LR (DTDR-L) (Ours) 27.8 0.75 0.63
Table 2: Retrieval performance for methods from different categories (QTS = Query-Tool Similarity, LR = Learned Retriever, DR = Dependency Retriever). QTS and LR categories are query-conditioned, while DR is history-conditioned. Our Dynamic methods are conditioned on both query and tool history. Function Selection Accuracy is reported on Qwen 3 0.6B using hard masking as ICL method. Best results per dataset and metrics are in bold. Using both query and history with DTDR retrieves relevant tools compared to only using one of them.
Dynamic Tool Dependency Retrieval-Linear (DTDR-L).
This supervised learning variant (see Figure 2, bottom-right) is based on a linear layer classifier trained to predict the set of next functions given the test query and the history of previous tool calls. We train a $1$-linear-layer classifier $\phi$ on top of a frozen embedding model $\theta$ to predict the next function given both the query and the function trajectory until $t$. Formally, $\phi(\theta(q+f_{0:t-1}))$, where the operator ++ denotes concatenation between strings. We use a softmax operation to obtain a probability distribution from the mode outputs. Let $P_{\phi}(f|q,f_{0:t-1}))$ be the probability of $f$ given the query and trajectory. To retrieve only a subset of tools, we set a threshold value $\alpha$, so that $\mathcal{F}_{t}=\omega(q,f_{0:t-1})=\{f|f\in\mathcal{F},P_{\phi}(f|q,f_{0:t-1})>\alpha\}$. The rows of the linear layer $\phi$ can be interpreted as learned embeddings of the tools $\mathcal{F}$, which are trained to align with the query embeddings $\theta(q)$. The number of learned parameters in this model is $e_{\theta}*|\mathcal{F}|$, only depending on the output size of the embedding model $e_{\theta}$ and number of tools $|\mathcal{F}|$. More implementation details in Appendix C.
5 Experimental Setup
Datasets.
We benchmark on four function calling datasets: TinyAgent (Erdogan et al. 2024), TaskBench DailyLife APIs, TaskBench HuggingFace, and TaskBench Multimedia (Shen et al. 2024). The first two datasets represent tools commonly available on a typical device, the other two datasets are specific to a particular domain, which makes them more challenging for “out-of-the-box” LLMs. We report in Appendix B the statistics for each benchmark, with an important insight being that all datasets have between 1 and 2 tool dependencies per plan, except for TaskBench DailyLife APIs which has almost zero dependencies. For each dataset, we set aside around 30% of the data for testing, while the reset is used as demonstration trajectories for tool retrieval and ICL methods.
LLM backbones.
We evaluate tool-retrieval methods and ICL strategies applied to seven LLM backbones. We consider the Qwen 3 family of models (0.6B, 1.7B, 4B, 8B, 14B) (Yang et al. 2025) representing edge devices of varying size, GPT-4o (Hurst et al. 2024) representing a cloud model solution, and Gorilla-V2 (Cheng-Jie Ji et al. 2024) representing an edge-device fine-tuned on function calling data. Assuming INT4 quantization and a KV cache of up to 10 thousand tokens, Qwen 3 models up to Qwen 3 4B could efficiently run on a typical mobile device (Federici et al. 2025; Song et al. 2025).
Figure 3: Comparison of efficient ICL methods against ICL with Raw Demonstrations and the baseline without ICL. All results are conditioned using the DTDR-C retriever. Pruning irrelevant tools from the prompt has a greater impact for smaller models that cannot handle longer contexts.
Metrics.
For downstream performance, we measure Function Selection Accuracy as described in (Erdogan et al. 2024): for each step, the selected function is considered correct only it belongs to the set of acceptable tools according to the Directed Acyclic Graph (DAG) of the ground-truth plan. For Success Rate we evaluate both function selection and predicted parameters for all steps in the plan. The success rate is 1 only if the DAG of the plan is isomorphic to the DAG of the ground-truth plan (Erdogan et al. 2024). The same LLM backbone is prompted to predict the function calling parameters given the query and selected function. Example of prompts for function selection and parameter filling are provided in Appendix F. Tool Retrieval methods are evaluated with ranking and retrieval metrics. Mean Reciprocal Rank (MRR) measures the relative ranking of tools. $\mathit{F_{1}}$ score is the harmonic mean of Precision and Recall and is computed considering the top-$k$ highest-scoring tools as retrieved, with $k$ being the number of acceptable tools according to the DAG of the ground-truth plan. We assess efficiency by: prompt length as a proxy for prefill speed (or time to encode and process the whole prompt); and number of parameters to quantify the LLM footprint which determines on-device usability.
Tool Retrievers Baselines.
We consider different categories of Tool Retrievals baselines from recent work presented in Table 1. Classical term-similarity baselines like BM-25 (Robertson, Zaragoza et al. 2009) directly compare text similarity between documents. Query-Tool Similarity (QTS) baselines use pre-trained sentence-embedding models to encode descriptions of queries and tools, without learning from demonstrations. This approach is extended in Less-is-More Level 1 (Paramanayakam et al. 2025) by prompting an LLM to describe the ideal tool set to solve a query, and in Tool Graph Retrieval (Gao et al. 2025) by additionally encoding tool dependencies in the embeddings. Learned Retriever (LR) baselines as in Erdogan et al. (2024) fine-tune small classifiers to learn what functions are useful to resolve a query. Finally, Dependency Retriever (DR) methods like ToolNet (Liu et al. 2024a) leverage tool dependencies to directly sub-select viable tools based on the most recent tool in the plan. DTDR introduces dynamic retrieval methods that leverage both the query and tool-call history to improve tool prediction. These are referred to as a Dynamic Learned Retriever (DTDR-L) and a Dynamic Dependency Retriever (DTDR-C). Additional implementation details for baselines and methods in Appendix C.
ICL Methods for Encoding $\mathcal{F}_{t}$ into Prompt.
For all methods including the No ICL baseline, we provide 1 randomly selected demonstration to show how to complete an example task. For all other ICL methods, we encode the subset of tools obtained from the Tool Retriever. For the Raw Demonstrations method, we provide up to $5$ additional demonstrations for which the plan includes the predicted tools. For a more efficient prompting, we can either use the Hard Mask and completely exclude the remaining tools from the prompt, or a Soft Mask, where the full set of tools is presented to the model but the retrieved list is emphasized as the set the model should generally prefer. If a Tool Retriever also predicts scores for each tool, these can be included in the prompt, yielding Weighted variants of the two masking approaches. Example of ICL prompts are provided in Appendix F.
6 Results and Discussion
Does DTDR improve tool retrieval?
We evaluate Tool Retrievers on ranking, retrieval and downstream performance (see Figure 2). MRR scores generally correlate well with Function Selection Accuracy, indicating that better ranking improves LLM reasoning. Set-based metrics like $F_{1}$ score are less informative of the downstream performance, as they do not capture the relative importance of tools. Overall, Learned Retriever and Dependency Retriever methods perform better than other baselines. Our dynamic Learned Retriever DTDR-L surpasses the static counterpart (Erdogan et al. 2024) by over 35% on TinyAgent and TaskBench DL, while matching performance on TaskBench HF and MM. This demonstrates the benefit of conditioning the prediction on the function selection history. Similarly, DTDR-C improves over the static Dependency Retriever counterpart (Liu et al. 2024a) by 50–100%, confirming the value of leveraging query and full history for retrieval, instead of only considering the latest function in the plan. Both BM-25 and Vanilla QTS retrievers reach lower performance. This is partially because they compare the query with tool descriptions, and since the “end-of-plan” function is never explicitly described in the queries, it is never part of the selected functions. In terms of efficiency, BM-25 and the DR baseline are model-free; other retrievers use a small sentence encoder to represent the query, whose cost is negligible relative to the subsequent LLM prompting step. The only exception is the retriever by Paramanayakam et al. (2025), which additionally prompts the LLM to describe the ideal tool set for the task.
Method TinyAgent TB-DL TB-HF TB-$M$
FSA ($\uparrow$) FCA ($\uparrow$) FSA ($\uparrow$) FCA ($\uparrow$) FSA ($\uparrow$) FCA ($\uparrow$) FSA ($\uparrow$) FCA ($\uparrow$)
Qwen3 0.6B No ICL 23.2 00.4 12.8 00.0 05.5 00.0 07.3 00.0
Raw Demonstrations 17.2 00.1 10.5 00.0 04.3 00.0 05.3 00.0
DR (Liu et al. 2024a) 22.0 00.2 11.5 00.0 06.7 00.0 09.5 00.3
LR (Erdogan et al. 2024) 19.8 00.9 19.6 02.0 10.0 00.5 12.1 00.7
DTDR-C (Ours) 35.3 04.2 23.7 06.8 20.8 05.9 20.2 08.3
DTDR-L (Ours) 46.1 03.5 45.9 00.8 18.9 01.2 17.6 03.7
Qwen3 1.7B No ICL 48.1 04.4 66.7 13.5 32.2 03.6 38.1 04.9
Raw Demonstrations 46.3 06.7 69.0 22.9 43.1 19.9 46.9 19.2
DR (Liu et al. 2024a) 51.6 02.6 60.0 09.7 41.6 05.7 45.1 06.8
LR (Erdogan et al. 2024) 50.5 07.1 56.6 20.3 42.8 17.3 42.1 17.3
DTDR-C (Ours) 64.1 11.8 60.2 26.4 52.8 21.3 47.8 22.5
DTDR-L (Ours) 78.4 14.5 83.1 29.4 49.0 19.7 45.5 20.4
Qwen3 4B No ICL 59.6 05.0 83.5 13.5 45.8 05.1 54.3 06.8
Raw Demonstrations 60.2 03.6 87.6 26.2 49.8 10.7 57.3 09.8
DR (Liu et al. 2024a) 62.7 09.2 64.7 19.3 52.8 19.3 51.8 16.8
LR (Erdogan et al. 2024) 52.2 09.0 60.4 19.3 56.9 28.0 58.2 26.5
DTDR-C (Ours) 68.1 14.4 64.7 27.9 56.0 35.1 52.4 28.0
DTDR-L (Ours) 80.7 16.9 89.0 31.8 60.5 29.8 64.1 30.9
Table 3: Comparison of selected methods in terms of function selection accuracy and end-to-end function calling accuracy over different datasets and edge-sized models. All retrievers are paired with weighted Hard Masking for ICL. Best results per dataset and model are bolded. Our DTDR reaches higher accuracy as it retrieves a set of tools more closely aligned with $\mathcal{F}_{t}^{*}$.
How to encode retrieved tools in the prompt?
In Figure 3, we analyze strategies to introduce retrieved tools or demonstrations in the LLM prompt. Adding demonstrations as raw text mainly benefits domain-specific datasets like TaskBench HF and MM, where names and descriptions of different tools are often similar. On three of four datasets, both Hard and Soft Masking outperform both the No ICL and the Raw Demonstration baselines. The exception is TaskBench DL, which lacks intrinsic function-call dependencies (see Table 4), making dependency-based prompting less relevant. Hard Masking directly prunes the set of available tools presented to the LLM, while Soft Masking only suggests most likely tools, leaving to the LLM the decision to pick from either these or from the full tool set. In other words, the choice between Hard and Soft masking corresponds to a trade-off between simplicity and expressiveness. For smaller model with limited capabilities, Hard Masking directly simplifies the problem and reaches best performance. On the contrary, larger models tend to exploit better the flexibility given by the Soft Masking approach. Except for Qwen 3 0.6B, Weighted Hard Masking performs on par with or better than its unweighted variant. Despite similar overall performance, we notice that weighted masking is better when assigning high probability to the correct tools, and unweighted masking is better in the opposite scenarios. This result suggests that unweighted masking should be preferred when expecting a large distribution shift between test queries and demonstration data, as predicted probabilities will be less accurate (details in Appendix E). Weighted Soft Masking follows the same trends and is omitted from the plot for clarity. From an efficiency standpoint, while Raw Demonstrations increase prompt length, Hard Masking reduces it substantially by removing descriptions of irrelevant tools.
How do DTDR perform on different datasets and model size?
In Table 3, we consider the best retrievers from previous experiments and evaluate their function selection accuracy (FSA) on several datasets and models of different sizes. More FSA results across different models and datasets can be found in Table 7 in the Appendix. DTDR-C consistently outperforms its static counterpart across all settings, while DTDR-L is comparable or better than the Learned Retrieval. This observation confirms that dynamic retrieval is key for high-quality function calling. Among all methods, DTDR-L achieves the best overall performance. Notably, for TinyAgent, DTDR-L applied to Qwen 3 4B and 8B respectively surpasses the No ICL baseline applied to Qwen 3 14B and GPT-4o, bridging the performance gap between edge and cloud models. The Raw Demonstrations approach hurts smaller models but become optimal at larger scales, where LLMs exhibit stronger reasoning capabilities.
Does DTDR improve end-to-end function calling success rates?
Also in Table 3, we report end-to-end Success Rate for the function calling task (FCA), using the same model with different prompts for the function selection and parameter filling subtasks (details in Appendix F). If one or more mistakes occurs at any step in the plan (either in the function name or its parameters), the whole plan is evaluated as unsuccessful. For datasets with tool dependencies (TA, TB-HF, TB-MM), the dynamic variants of DTDR improve the end-to-end Success Rate between 300% and 600% compared to the baseline without ICL, and between 23% and 104% compared to the best Tool Retrieval baseline. Weaker baselines on smaller models fail on all datasets. The Learned Retrieval (Erdogan et al. 2024) tends to overfit to predicting the most common function in the dataset, while the same architecture in its dynamic variant DTDR-L (adding function call history as input), prevents this problem and yields significantly better performance. The baseline using Raw Demonstrations in the prompt reaches good performance with larger models, which have increased capabilities of reasoning on long text. However, as we discuss next, using long demonstrations in the prompt would make the system latency significantly higher.
Figure 4: Prompt length across different methods and datasets. Our method reduces the prompt length by: 1) efficiently encoding ICL examples as tool dependencies instead of Raw Demonstrations, and 2) only retrieving tool dependencies which are relevant for the test query.
Figure 5: Ablations on: a) history length, b) # of k-means clusters, and c) # of demonstrations.
What is the prompt efficiency of different retrievers?
In Figure 4, we analyze the same methods from a prompt-efficiency perspective. Since function selection requires generating only a few tokens, inference time is dominated by the prefill step. We therefore compare prompt lengths across methods and datasets, distinguishing between constant and variable prompt sections (see examples in Appendix F). The constant section (guidelines, task demonstrations, and the tool descriptions) can be precomputed and cached, while the variable section (query encoding and ICL-specific text) must be recomputed for each query. For the No ICL baseline the function descriptions occupy most of the prompt, which further increases in size when adding Raw Demonstrations. This quickly scales up as the number of tools increases (e.g. TaskBench HF). Static Dependency Retrieval reduces this by selecting a subset of functions, but the subset changes at each step and must be updated dynamically in the prompt. While the Learned Retriever can select a narrow set of tools, it tends to overfit to retrieving the most frequent tool, resulting in low end-to-end performance. Our DTDR variants go further: dynamic retrieval narrows the candidate set of tools significantly, reducing total prompt length by up to 51% and the variable portion by up to 72%. Combined with accuracy gains, these results highlight the dual benefit of DTDR: improved reasoning and enhanced prompt efficiency.
What is the impact of History Length, Clustering, and Number of Demonstrations?
Figure 5a examines the effect of varying the history length $l$ in the function call sequence $f_{t-l},...,f_{t-1}]$. Longer histories enhance dependency modeling in DTDR-C and provide richer context for DTDR-L. Both models benefit up to $l=3$, beyond which gains taper off. Without history, DTDR-L tends to overfit, often defaulting to the most frequent functions, as previously observed. Figure 5b explores query-conditioning by varying the number of clusters in DTDR-C. With a single cluster, retrieval is query-agnostic, similar to (Liu et al. 2024a). Increasing the number of clusters enables more targeted retrieval from demonstrations similar to the test query. The extreme case (nearest-neighbor retrieval) assigns one demonstration per cluster. We find optimal performance when the number of clusters is approximately $1/10$ of the demonstrations, which we adopt as the default setting for all DTDR-C experiments. Figure 5c shows how increasing the number of demonstrations influences Function Selection Accuracy for both DTDR-C and DTDR-L. Performance improves with more data, plateauing around 10k samples. Gains are more pronounced for smaller LLM backbones. Notably, DTDR-L suffers a sharp drop below 1k demonstrations due to overfitting, highlighting its sensitivity compared to the more robust DTDR-C.
7 Conclusion and Future Work
We introduce Dynamic Tool Dependency Retrieval (DTDR), a method for task-aligned retrieval that captures multi-step tool dependencies. DTDR leverages both the query and historical predictions via a retrieval module trained on demonstrations, with two variants: a supervised approach and an unsupervised clustering-based approach, both suitable for low-resource, on-device settings. Experiments across multiple datasets and LLM backbones show DTDR significantly improves retrieval quality and downstream function selection over baselines. We further analyze prompt efficiency, history context, demonstration count, and clustering effects. Future directions include handling imperfect demonstrations, extending to multimodal tool-based tasks (e.g. robotics), and adapting to evolving tool sets.
Appendix A Detailed Related Works
Training LLMs for Tool Use.
Recent LLMs demonstrate impressive tool-usage capabilities, with cloud-based models such as GPT-5 (OpenAI 2025), Claude 4 (Anthropic 2025), and GLM-4.5 (Zeng et al. 2025) leading by a wide margin on function-call benchmarks like BFCL V4 (Patil et al. 2025) and $\tau^{2}$-Bench (Barres et al. 2025). In contrast, smaller LLMs designed for edge deployment perform significantly worse on agentic tasks when evaluated out-of-the-box, creating a challenge for widespread adoption on mobile devices. To address this, recent work has explored fine-tuning LLMs on large function calling datasets, sourced from cloud providers (Cheng-Jie Ji et al. 2024), generated by cloud models (Qiao et al. 2023), or obtained through self-play in agentic environments (Schick et al. 2023). However, collecting tool-use data for fine-tuning is often prohibitively expensive, and these fine-tuned models frequently struggle to generalize to new or updated tools.
LLM Prompting with Tool Retrieval.
An alternative line of research leverages the in-context learning capabilities of LLMs, introducing prompt-engineering strategies for tool learning. This can involve including tool descriptions (Shen et al. 2023, 2024) or example trajectories (Paranjape et al. 2023; Sarukkai, Xie, and Fatahalian 2025) within the model prompt. To handle hundreds of apps and APIs, recent work uses retrieval-augmented generation to sub-select tools that are relevant to the task. ToolLLM (Qin et al. 2023) and ToolReAGt (Braunschweiler, Doddipatla, and Zorila 2025) identify the most suitable tools by comparing embeddings derived from the user query and tool documentation. Less-is-More (Paramanayakam et al. 2025) further prompts an LLM to generate documentation for the ideal tools to solve the query. ART (Paranjape et al. 2023) uses LLM prompting to retrieve example trajectories for most relevant tools.
Tool Dependency Retrieval.
The execution of certain tools often depends on the outputs of others. ControlLLM (Liu et al. 2024b) and SciToolAgent (Ding et al. 2025) address specific scenarios where the tool dependency graph is known.In most real-world cases, however, the set of tools and their dependencies are not known in advance or may change at test time, for example when apps are installed or updated. In that case, tool dependencies could be learned from demonstration data provided by tool vendors (Paranjape et al. 2023), extracted from historical tool trajectories (Chen et al. 2025), generated procedurally (Erdogan et al. 2024; Shen et al. 2024), or synthesized using cloud LLMs (Gao et al. 2025; Patil et al. 2024; Qin et al. 2023). Tool Graph Retriever (Gao et al. 2025) and GTool (Chen et al. 2025) employ GNNs to encode the learned dependencies, using the resulting embeddings for similarity-based retrieval and for LLM prompting. In contrast, ToolNet (Liu et al. 2024a) parses the tool dependency graph to identify the set of eligible tools after the latest tool call in the plan. Rather than modeling a static tool graph from demonstration data, we propose a lightweight approach that dynamically retrieves specific tool dependencies which are relevant to the test query and the sequence of previously called tools.
Appendix B Dataset Statistics
In Table 4 we report important statistics for the datasets considered in our evaluation. While the average length of the plans is comparable across datasets, the number of tool dependencies varies significantly, which means that certain datasets benefit more than others in modeling and employing tool dependencies for function selection.
Dataset # Plans # Tools # Tool calls per plan # Tool dependencies per plan
[…]
Table 4: Function Calling Benchmark statistics. The number of tool dependencies refers to cases where the parameter of a function call is the output of a function call earlier in the plan. Notice that TaskBench DailyLife mainly contains plans with unrelated function calls.
Tool Dependency Graph Unweighted Tool, Weighted Tool
$G_{6}$ Key, $f_{6}$ Dependency Graph Value, $F^{f_{6}}$ Dependency Graph Value, $P(f^{\prime}|f_{6})$
(‘start’, [‘join’, {‘join : $0.8$,
‘start’, ‘create_calendar_event’, ‘create_calendar_event’ : $0.05$,
‘start’, ‘create_reminder’, ‘create_reminder’ : $0.05$,
‘open_and_get_file_path’, ‘summarize_pdf’, ‘summarize_pdf’: $0.05$,
‘get_email_address’, ‘append_note_content’] ‘append_note_content’: $0.05$ }
‘compose_new_email’)
Table 5: Example Tool Dependency Graph. Given a function sequence of length $6$, $f_{6}$ (left), we can compute based on demonstration data, $D$, the unweighted tool dependency graph (middle) and weighted tool dependency graph (right).
Function Selection Acc. [%] ($\uparrow$)
Retrieval Metrics
Retrieval Method   MRR ($\uparrow$)   F1 ($\uparrow$)
TinyAgent Random Guess 05.9 0.26 0.12
BM-25 (Robertson, Zaragoza et al. 2009) 23.1 0.35 0.20
QTS (Vanilla) 15.8 0.26 0.14
QTS (Gao et al. 2025) 21.5 0.18 0.09
QTS (Paramanayakam et al. 2025) 23.7 0.36 0.19
DR (Liu et al. 2024a) 30.7 0.70 0.49
Dynamic DR (DTDR-C) (Ours) 43.3 0.78 0.56
LR (Erdogan et al. 2024) 25.6 0.53 0.39
Dynamic LR (DTDR-L) (Ours) 65.1 0.93 0.55
Taskbench-HF Random Guess 04.2 0.16 0.05
BM-25 (Robertson, Zaragoza et al. 2009) 08.3 0.18 0.05
QTS (Vanilla) 07.9 0.21 0.08
QTS (Gao et al. 2025) 14.6 0.37 0.27
QTS (Paramanayakam et al. 2025) 12.9 0.30 0.18
DR (Liu et al. 2024a) 17.1 0.46 0.33
Dynamic DR (DTDR-C) (Ours) 27.5 0.64 0.55
LR (Erdogan et al. 2024) 20.5 0.53 0.32
Dynamic LR (DTDR-L) (Ours) 27.8 0.75 0.63
Taskbench-$M$ Random Guess 02.4 0.11 0.03
BM-25 (Robertson, Zaragoza et al. 2009) 13.7 0.26 0.19
QTS (Vanilla) 05.3 0.14 0.04
QTS (Gao et al. 2025) 05.3 0.14 0.20
QTS (Paramanayakam et al. 2025) 07.6 0.29 0.18
DR (Liu et al. 2024a) 13.3 0.38 0.27
Dynamic DR (DTDR-C) (Ours) 24.4 0.52 0.43
LR (Erdogan et al. 2024) 21.0 0.49 0.28
Dynamic LR (DTDR-L) (Ours) 27.0 0.69 0.55
Taskbench-DL Random Guess 02.4 0.15 0.07
BM-25 (Robertson, Zaragoza et al. 2009) 13.5 0.31 0.15
QTS (Vanilla) 07.5 0.16 0.07
QTS (Gao et al. 2025) 13.3 0.17 0.06
QTS (Paramanayakam et al. 2025) 18.9 0.40 0.26
DR (Liu et al. 2024a) 14.7 0.36 0.18
Dynamic DR (DTDR-C) (Ours) 28.2 0.54 0.29
LR (Erdogan et al. 2024) 23.6 0.55 0.44
Dynamic LR (DTDR-L) (Ours) 58.9 0.85 0.64
Table 6: Retrieval performance for methods from different categories (QTS = Query-Tool Similarity, LR = Learned Retriever, DR = Dependency Retriever). QTS and LR categories are query-conditioned, while DR is history-conditioned. Our Dynamic methods are conditioned on both query and tool history. Function Selection Accuracy is reported on Qwen 3 0.6B using hard masking as ICL method. Best results per dataset and metrics are in bold.
Appendix C Implementation Details
Tool Dependency Graph Construction in DTDR-C
We compute the tool dependency graph based on the training data $D_{train}$. Each training datapoint contains a ground-truth trajectory for the given query. Let $f_{l}=(f_{i-o},f_{i-l+1},\cdots,f_{i})$ be a history of function calls of length $l$. For example, a possible $f_{2}$ could be the tuple, (’get_email’, ’reply_email’), where ‘get_email’ is the first function in the sequence and ‘reply_email is the latest function. For each function sequence $f_{l}$ used in the ground-truth trajectories of $D_{train}$, let $\mathcal{F}^{f_{l}}$ be the set of all unique functions that have appeared immediately after $f_{l}$. The tool dependency graph $G_{l}$ is formatted according to whether it is probabilistic or not. When $G_{l}$ is probabilistic, each $f_{l}$ is a key in a nested dictionary. The value of each $f_{l}$ is another dictionary representing a probability distribution where the inner keys are all the functions in $\mathcal{F}^{f_{l}}$. For each inner function key, $f^{\prime}\in\mathcal{F}^{f_{l}}$, the value is, $P(f^{\prime}|f_{l})$: the percent of times $f^{\prime}$ appears given that $f_{l}$ appears immediately before it. Thus, $P(f^{\prime}|f_{l})$ can be considered an $l$-ordered Markov Chain. If $G_{l}$ is unweighted, then $G_{l}$ is a flat dictionary where each $f_{l}$ is a key and the corresponding value is $\mathcal{F}^{f_{l}}$. Table 5 gives and example of a tool dependency graph of order $6$, or $G_{6}$. To pad function sequences with length less than order $l$, we use the string ‘start’ as a placeholder. In Table 5, the example $f_{6}$ only has three functions in the history, so we use use three ‘start’ strings in the beginning. In summary, we use a sliding window approach with window length $l$, and we construct $G$ based on the next function right after the current window.
Classifier optimization in DTDR-L
For DTDR-C, we find the optimal number of clusters to be $1/10$ of the training set size. Fine-tuning for TinyAgent and DTDR-L variants is performed using Adam with a learning rate of $10^{-3}$, learning rate decay of $0.9$, weight decay of $10^{-5}$, and $10$ epochs. For DTDR-L, the training of $\phi$ is per-step. For each training query $q\in Q_{train}$, we train our predictor $\omega$ using the sum of independent $|\mathcal{F}|$ binary cross-entropy loss terms. Formally, let $d_{0:t-1}$ be the possible ground-truth demonstration trajectory for $q$ up until time $t$. Both $d_{0:t-1}$ and $f_{0:t-1}$ are a list of function names, so $d_{0:t-1}$ can be concatenated to $q$ and passed into $\phi$. Therefore, we can write
$\begin{split}\phi^{*}&=\\ &\operatorname*{arg\,min}_{\phi}-\sum_{q\in Q_{train}}\sum_{t=0}^{T}\sum_{f\in\mathcal{F}}(\mathbf{1}_{f\in\mathcal{F}_{q,t}^{*}})(\log\phi(q,d_{0:t-1})).\end{split}$ (2)
Other details and hyper-parameters
For all Tool Retriever methods, we use the same sentence-embedding model paraphrase-MiniLM-L6-v2 (Reimers and Gurevych 2019) across all approaches to ensure comparability. For Less-is-More, we ensure the same LLM is used both for Function Selection and for recommending the ideal tool descriptions. We consider Search Level 1 but not Level 2, as the latter needs interaction with a cloud model, which is prohibitive on-device. For ToolNet, we build the tool dependency graph from the training split and use it to retrieve and score relevant tools based on the latest tool call in the plan. For a fair comparison with Tool Graph Retrieval, we also use the available example trajectories to build the tool dependency graph, instead of relying on a separate predictor. For the TinyAgent retriever, we fine-tune a linear layer on top of the sentence embedding model to score all tools given the test query. For all Tool Retrievers other than DTDR-C (which does not require it) we sweep the function selection threshold with values of $0.2,0.23,0.5$ and set it based on function selection accuracy. For ICL experiments we use greedy decoding with a maximum generation length matching the longest tool name in the dataset. For the Raw Demonstrations ICL method, we provide 5 random demonstrations as examples. When defining train and test splits for each dataset, we ensure that the number of occurrences in train and test data for each tool follows then 70/30 ratio. We use FP16 precision for Qwen models and Gorilla-V2, while GPT-4o is prompted through cloud APIs. All experiments are conducted on NVIDIA A100 GPUs. We set the seed value to $5006$ for all experiments.
Appendix D Additional Results
Retrieval Accuracy
We report in Table 6 the full values for retrieval and function selection accuracy with Qwen 3 0.6B with unweighted Hard Masking. DTDR consistently outperforms in both retireval and function selection across datasets. We note that for Taskbench-Multimedia that LR outperforms in function selection due to overfitting to repeatedly predict “end-of-plan” which is always the last function in each task.
Function Selection Accuracy Across Models
We provide in Table 7 extended results on function selection accuracy. Additional LLM backbones include Qwen 3 8B, Qwen 3 14B, Gorilla-v2 and GPT-4o. At similar footprints, Gorilla V2 under-performs Qwen 3 8B, likely due to differences between our test data and the coding API data used for Gorilla’s fine-tuning. This difference in performance reinforces our claim that adaptive ICL strategies are preferable to fine-tuned generalist models, as ICL can seamlessly adapt to unseen tools at test time.
Function Selection Accuracy [%] ($\uparrow$)
Method TinyAgent TaskBench DailyLife TaskBench HuggingFace TaskBench Multimedia
Qwen3 0.6B No ICL 23.2 12.8 05.5 07.3
Raw Demonstrations 17.2 10.5 04.3 05.3
Dependency Retrieval (Liu et al. 2024a) 22.0 11.5 06.7 09.5
Learned Retrieval (Erdogan et al. 2024) 19.8 19.6 10.0 12.1
DTDR-C (Ours) 35.3 23.7 20.8 20.2
DTDR-L (Ours) 46.1 45.9 18.9 17.6
Qwen3 1.7B No ICL 48.1 66.7 31.2 38.1
Raw Demonstrations 46.3 69.0 43.1 46.9
Dependency Retrieval (Liu et al. 2024a) 51.6 60.0 41.6 45.1
Learned Retrieval (Erdogan et al. 2024) 50.5 56.6 42.8 42.1
DTDR-C (Ours) 64.1 60.2 52.8 47.8
DTDR-L (Ours) 78.4 83.1 49.0 45.5
Qwen3 4B No ICL 59.6 83.5 45.8 54.3
Raw Demonstrations 60.2 87.6 49.8 57.3
Dependency Retrieval (Liu et al. 2024a) 62.7 64.7 52.8 51.8
Learned Retrieval (Erdogan et al. 2024) 52.2 60.4 56.9 58.2
DTDR-C (Ours) 68.1 64.7 56.0 52.4
DTDR-L (Ours) 80.7 89.0 60.5 64.1
Qwen3 8B No ICL 71.3 93.9 70.5 78.5
Raw Demonstrations 67.8 94.5 72.5 80.4
Dependency Retrieval (Liu et al. 2024a) 76.7 75.3 67.3 69.1
Learned Retrieval (Erdogan et al. 2024) 43.7 52.2 52.0 49.6
DTDR-C (Ours) 80.4 75.3 67.3 69.1
DTDR-L (Ours) 89.0 91.0 63.8 67.6
Qwen3 14B No ICL 78.0 96.2 74.4 81.7
Raw Demonstrations 77.9 96.6 76.7 82.3
Dependency Retrieval (Liu et al. 2024a) 79.5 84.1 69.2 73.8
Learned Retrieval (Erdogan et al. 2024) 51.1 56.5 53.5 50.8
DTDR-C (Ours) 84.8 84.1 71.5 74.0
DTDR-L (Ours) 92.5 93.7 71.2 73.4
Gorilla V2 No ICL 39.5 PLE PLE PLE
Raw Demonstrations PLE PLE PLE PLE
Dependency Retrieval (Liu et al. 2024a) 32.2 PLE PLE PLE
Learned Retrieval (Erdogan et al. 2024) 37.9 PLE PLE PLE
DTDR-C (Ours) 43.4 PLE PLE PLE
DTDR-L (Ours) 58.7 PLE PLE PLE
GPT-4o No ICL 84.7 96.6 74.8 81.6
Raw Demonstrations 86.8 96.8 77.1 82.3
Dependency Retrieval (Liu et al. 2024a) 85.9 82.9 70.0 72.8
Learned Retrieval (Erdogan et al. 2024) 53.2 48.1 55.2 51.1
DTDR-C (Ours) 85.9 82.9 71.9 74.8
DTDR-L (Ours) 94.2 91.5 71.8 73.7
Table 7: Comparison of selected methods in terms of function selection accuracy over different datasets and models. Best results per dataset and model are bolded. Gorilla V2 has 4096 maximum context length, which easily results in PLE (Prompt Length Exceeded).
Appendix E Weighted vs Unweighted Masking
To see the impact between weighted and unweighted masking, we compute the percentage of correct predictions when the DTDRCDTDR-C-based $\omega$ did not retrieve any $f\in F^{*}_{t}$ with a probability above threshold $0.5$. We see for Tinyagent across all models and number of cluster settings, Hard Masking unweighted achieves the maximum of around $33.6\%$ and outperforms hard mask weighted by up to $10\%$. This difference is also present when comparing Soft Masking with its weighted variant, and this difference shows that when the demonstration data does not reflect the task well enough, the weights provided in the prompt mislead the agent. However, as the number of clusters increase, this difference between in values for weighted and unweighted decreases. Similar observations are present for probability thresholds of $0.4$ and $0.3$.
Appendix F Prompt Examples
We provide example prompts showing guidelines and ICL techniques used for function selection and parameter filling. Grayed out comments are only used as reference, and are not included in the actual prompt.
Full Prompt F.1: Function Selection Full Prompt F.2: Parameter Filling ICL Prompt F.1: No ICL ICL Prompt F.2: Soft Mask ICL Prompt F.3: Hard Mask ICL Prompt F.4: Hard Mask (Weighted)
Appendix G Use of Large Language Models
Microsoft Co-Pilot was used for searching for related works and providing edit suggestions to the manuscript.